# Light-induced odd-parity altermagnets on dimerized lattices

### v1

@@ Abstract
For convenience, we have introduced the following shorthand notations: $\tilde{t}_{x}=t_{+}J_{0}(A_{1})$, $\tilde{t}_{y}=2t_{2}J_{0}(A_{2})$, $\tilde{t}_{z}=2t_{3}J_{0}(A_{3})$, $F_{y}=8J_{1}(A_{1})J_{1}(A_{2})t_{-}t_{2}\sin(\phi_{2}-\phi_{1})/\omega$, $F_{z}=8J_{1}(A_{1})J_{1}(A_{3})t_{-}t_{3}\sin\phi_{1}/\omega$, where $A_{i}=\lambda_{i}A_{0}$, $\lambda_{1}=\sqrt{1-\cos^{2}\phi\sin^{2}\theta}$, $\lambda_{2}=\sqrt{1-\sin^{2}\phi\sin^{2}\theta}$, $\lambda_{3}=\eta\sin\theta$, $\phi_{1}=\arg[\eta\cos\phi\cos\theta+i\sin\phi]$, and $\phi_{2}=\arg[\eta\sin\phi\cos\theta-i\cos\phi]$. Despite the added complexity introduced by general-angle incidence, the resulting spin-splitting features and band topology remain straightforward to analyze.
@@
Given the apparent symmetry between $k_{y}$ and $k_{z}$ in the Hamiltonian, Fermi surfaces with opposite spins can now be shifted in opposite directions along any axis in the $k_{y}$-$k_{z}$ plane. To illustrate this symmetry, we plot the spin-split Fermi surfaces under CPL incident along the $y$ and $z$ directions. As shown in Fig. 4(b), Fermi surfaces with opposite spins exhibit opposite displacements: along the $z$ direction for $y$-propagating CPL, and along the $y$ direction for $z$-propagating CPL. Notably, when CPL propagates along any direction in the $yz$ plane, the dimerization axis, the light propagation direction, and the Fermi surface displacement direction form an approximately orthogonal triad.
@@
Figure 4: (a) Realization of the 3D system through AB stacking. (b) Fermi surfaces at $E_{F}=0.8$ for spin up (red) and spin down (blue), under CPL incident along the $z$ (left) and $y$ (right) directions. (c) Energy spectrum in the $k_{x}=0$ plane, showing two Weyl points of each spin. (d) Berry curvature vector field across the Brillouin zone. The arrows’ direction and magnitude represent the local Berry curvature. The value $M=1$ is used in (b), and $M=0.4$ is used in (c) and (d). Other shared parameters are $t_{0}=1$, $t_{1}=0.2$, $t_{2}=0.5$, $t_{3}=2$, $\eta=1$, $A_{0}=1.8$, and $\omega=9$.
@@
In our 2D system, we demonstrate that a Chern insulator emerges when the CPL amplitude surpasses a critical threshold. Stacking such Chern insulators can produce either 3D Chern insulators (for weak interlayer coupling) or Weyl semimetals (for strong coupling). Here, we focus on the realization of Weyl semimetals. From the 3D Hamiltonian [Eq. (8)], Weyl points emerge in the $k_{x}=0$ plane when the CPL amplitude surpasses another critical threshold. Their positions are determined by: $\tilde{t}_{x}+\tilde{t}_{y}\cos k_{y}+\tilde{t}_{z}\cos k_{z}=0$ and $sM+F_{y}\sin k_{y}+F_{z}\sin k_{z}=0$ ($s=\pm 1$ for spin up/down). This system contains a minimum of four Weyl points, mirroring the situation in noncentrosymmetric Weyl semimetals with TRS [93]. Notably, although the unitary $C_{2\perp}\|\mathcal{P}]$ symmetry constrains Weyl point locations identically to antiunitary TRS, it imposes opposite rules on the monopole charges of symmetry-related points. Specifically, Weyl points related by $C_{2\perp}\|\mathcal{P}]$ carry opposite charges, whereas those related by TRS carry the same charge.
@@
To demonstrate our analysis, we consider the simplified case of $y$-direction ($\phi=\theta=\pi/2$) right-handed ($\eta=1$) CPL incidence for clarity. In this configuration, the possible spin-up Weyl points are located at $\boldsymbol{k}_{w1}=(0,k_{y}^{+},k_{z}^{w})$, $\boldsymbol{k}_{w2}=(0,-k_{y}^{+},k_{z}^{w})$, $\boldsymbol{k}_{w3}=(0,k_{y}^{-},\pi-k_{z}^{w})$, $\boldsymbol{k}_{w4}=(0,-k_{y}^{-},\pi-k_{z}^{w})$, where $k_{z}^{w}=-\arcsin(M/F_{z})$, and $k_{y}^{\pm}=\pi-\arccos[(\tilde{t}_{x}\pm\tilde{t}_{z}\cos k_{w}^{z})/\tilde{t}_{y}]$. By choosing the parameters such that $|\tilde{t}_{x}-\tilde{t}_{z}\cos k_{w}^{z}|<\tilde{t}_{y}<|\tilde{t}_{x}+\tilde{t}_{z}\cos k_{w}^{z}|$, only the two Weyl points at $\boldsymbol{k}_{w3}$ and $\boldsymbol{k}_{w4}$ are present. Correspondingly, there are two spin-down Weyl points located at $\boldsymbol{k}_{w3}$ and $\boldsymbol{k}_{w4}$. In Fig. 4(c), we plot the energy bands in the $k_{x}=0$ plane, which clearly exhibit two Weyl cones in each spin sector. Furthermore, Fig. 4(d) shows the distribution of Berry curvature vector field $\boldsymbol{\Omega}(\boldsymbol{k})=(\Omega_{x}(\boldsymbol{k}),\Omega_{y}(\boldsymbol{k}),\Omega_{z}(\boldsymbol{k}))$, where $\Omega_{a}(\boldsymbol{k})=\sum_{E_{n}<0}i\epsilon_{abc}\langle\partial_{k_{b}}u_{n}|\partial_{k_{c}}u_{n}\rangle$. Here $\epsilon_{abc}$ is the Levi-Civita symbol and $|u_{n}(\boldsymbol{k})\rangle$ is the eigenstate of the $n$-th occupied band. The field clearly reveals that inversion-related Weyl points act as a source and a sink of Berry flux, confirming they carry opposite monopole charges [93].
Discussions and conclusions.—We demonstrate that CPL can dynamically induce odd-parity altermagnetism in collinear $\mathcal{PT}$-symmetric antiferromagnets on dimerized lattices. The resulting odd-parity altermagnetic phase inherently supports topological states due to the underlying Dirac band structure. Remarkably, these systems also host favorable conditions for the emergence of unconventional superconductivity, as their band structure mimics time-reversal-invariant systems with SOC while explicitly breaking TRS. The fundamental relation $E(\mathbf{k},\boldsymbol{s})=E(-\mathbf{k},-\boldsymbol{s})$ ensures that a spin-up electron always has a spin-down partner available for pairing. This opens the door to realizing exotic, time-reversal-breaking spin-singlet superconductors.
@@
Antiferromagnets on dimerized lattices correspond to systems exhibiting coexisting bond order and antiferromagnetic order. The bond order typically arises from electron-phonon interactions, whereas the antiferromagnetic order is driven by strong electron-electron repulsion. Their coexistence has been rigorously established through numerical studies of strongly correlated systems [89]. Beyond correlated materials, cold-atom systems offer a promising alternative platform for realizing our proposed odd-parity AMs. In such systems, the dimerized lattice can be directly engineered through optical lattice design [94], while the antiferromagnetic order can be stabilized by tuning the on-site interactions to be strongly repulsive [95, 96, 97]. The effect of CPL can be equivalently realized by periodically modulating the lattice [98].
@@
To conclude, our work establishes a general dynamical protocol for creating odd-parity AMs, significantly expanding the class of systems capable of hosting these exotic phases and their emergent phenomena.

### v2

@@ Abstract
For convenience, we have introduced the following shorthand notations: $\tilde{t}_{x}={\color[rgb]{0,0,0}t_{+}}J_{0}(A_{1})$, $\tilde{t}_{y}=2t_{2}J_{0}(A_{2})$, $\tilde{t}_{z}=2t_{3}J_{0}(A_{3})$, $F_{y}=8J_{1}(A_{1})J_{1}(A_{2}){\color[rgb]{0,0,0}t_{-}}t_{2}\sin(\phi_{2}-\phi_{1})/\omega$, $F_{z}=8J_{1}(A_{1})J_{1}(A_{3}){\color[rgb]{0,0,0}t_{-}}t_{3}\sin\phi_{1}/\omega$, where $A_{i}=\lambda_{i}A_{0}$, $\lambda_{1}=\sqrt{1-\cos^{2}\phi\sin^{2}\theta}$, $\lambda_{2}=\sqrt{1-\sin^{2}\phi\sin^{2}\theta}$, $\lambda_{3}=\eta\sin\theta$, $\phi_{1}=\arg[\eta\cos\phi\cos\theta+i\sin\phi]$, and $\phi_{2}=\arg[\eta\sin\phi\cos\theta-i\cos\phi]$. Despite the added complexity introduced by general-angle incidence, the resulting spin-splitting features and band topology remain straightforward to analyze.
@@
Given the apparent symmetry between $k_{y}$ and $k_{z}$ in the Hamiltonian, Fermi surfaces with opposite spins can now be shifted in opposite directions along any axis in the $k_{y}$-$k_{z}$ plane. To illustrate this symmetry, we plot the spin-split Fermi surfaces under CPL incident along the $y$ and $z$ directions. As shown in Fig. 4(b), Fermi surfaces with opposite spins exhibit opposite displacements: along the $z$ direction for $y$-propagating CPL (left panel), and along the $y$ direction for $z$-propagating CPL (right panel). For CPL propagating along a generic direction in the $yz$ plane, defined by $(\phi,\theta)=(\pi/2,\theta)$ or $\boldsymbol{n}=(0,\sin\theta,\cos\theta)$, the parameters of the Floquet Hamiltonian are: $\phi_{1}=\pi/2$, $\phi_{2}=[1+\eta\text{sgn}(\cos\theta)]\pi/2$, $\lambda_{1}=1$, $\lambda_{2}=|\cos\theta|$, and $\lambda_{3}=\eta\sin\theta$. Consequently, we have $F_{y}=-8\eta\text{sgn}(\cos\theta)J_{1}(A_{0})J_{1}(A_{0}|\cos\theta|)t_{-}t_{2}/\omega$, and $F_{z}=8J_{1}(A_{0})J_{1}(\eta A_{0}\sin\theta)t_{-}t_{3}/\omega$. In the case where the $y$ and $z$ directions are equivalent ($t_{2}=t_{3}$), one gets the ratio $F_{y}/F_{z}=-\eta\text{sgn}(\cos\theta)J_{1}(A_{0}|\cos\theta|)/J_{1}(\eta A_{0}\sin\theta)=-J_{1}(A_{0}\cos\theta)/J_{1}(A_{0}\sin\theta)$. Now, considering the weak-driving regime ($A_{0}\ll 1$) and momenta near $(k_{y},k_{z})=(0,0)$, we can use the approximations $J_{1}(x)\approx x/2$ and $\sin k\approx k$, The ratio becomes $F_{y}/F_{z}\approx-\cos\theta/\sin\theta$, and the key term in the Hamiltonian has the property: $F_{y}\sin k_{y}+F_{z}\sin k_{z}\propto-\cos\theta k_{y}+\sin\theta k_{z}$. This term vanishes along the direction $(k_{y},k_{z})\parallel(\sin\theta,\cos\theta)$ and is maximized along the perpendicular direction $(k_{y},k_{z})\parallel(-\cos\theta,\sin\theta)$. Consequently, Fermi surfaces with opposite spins are displaced oppositely along $(-\cos\theta,\sin\theta)$, which is perpendicular to the light’s in-plane propagation direction $(\sin\theta,\cos\theta)$. This analysis reveals that when light propagates perpendicular to the dimerization axis, the three vectors—the dimerization axis, the light’s propagation direction, and the Fermi surface displacement—form an approximately orthogonal triad.
@@
Figure 4: (a) Realization of the 3D system through AB stacking. (b) Fermi surfaces of the CPL-driven system at a Fermi energy $E_{F}=0.8$. The red and blue contours represent the spin-up and spin-down branches, respectively. Left panel: CPL incident along the $z$ direction. Right panel: CPL incident along the $y$ direction. (c) Energy spectrum in the $k_{x}=0$ plane, showing two Weyl points of each spin. (d) Berry curvature vector field in the $k_{x}=0$ plane containing the Weyl points. Arrows show the in-plane components $(\Omega_{y},\Omega_{z})$. Red and blue dots mark the Weyl points, which act as Berry flux sources (arrows flowing out) and sinks (arrows flowing in), respectively. The value $M=1$ is used in (b), and $M=0.4$ is used in (c) and (d). Other shared parameters are $t_{0}=1$, $t_{1}=0.2$, $t_{2}=0.5$, $t_{3}=2$, $\eta=1$, $A_{0}=1.8$, and $\omega=9$.
@@
In our 2D system, we demonstrate that a $p$-wave altermagnetic Chern insulator emerges when the CPL amplitude surpasses a critical threshold. Stacking such Chern insulators can produce either 3D $p$-wave altermagnetic Chern insulators (for weak interlayer coupling) or Weyl semimetals (for strong coupling) [101, 102]. Here, we focus on the realization of $p$-wave altermagnetic Weyl semimetals. From the 3D Hamiltonian [Eq. (8)], Weyl points emerge in the $k_{x}=0$ plane when the CPL amplitude surpasses another critical threshold. Their positions are determined by: $\tilde{t}_{x}+\tilde{t}_{y}\cos k_{y}+\tilde{t}_{z}\cos k_{z}=0$ and $sM+F_{y}\sin k_{y}+F_{z}\sin k_{z}=0$ ($s=\pm 1$ for spin up/down). This system contains a minimum of four Weyl points, mirroring the situation in noncentrosymmetric Weyl semimetals with TRS [103]. Notably, although the unitary $C_{2\perp}\|\mathcal{P}]$ symmetry constrains Weyl point locations identically to antiunitary TRS, it imposes opposite rules on the monopole charges of symmetry-related points. Specifically, Weyl points related by $C_{2\perp}\|\mathcal{P}]$ carry opposite charges, whereas those related by TRS carry the same charge.
@@
To demonstrate our analysis, we consider the simplified case of $y$-direction ($\phi=\theta=\pi/2$) right-handed ($\eta=1$) CPL incidence for clarity. In this configuration, the possible spin-up Weyl points are located at $\boldsymbol{k}_{w1}=(0,k_{y}^{+},k_{z}^{w})$, $\boldsymbol{k}_{w2}=(0,-k_{y}^{+},k_{z}^{w})$, $\boldsymbol{k}_{w3}=(0,k_{y}^{-},\pi-k_{z}^{w})$, $\boldsymbol{k}_{w4}=(0,-k_{y}^{-},\pi-k_{z}^{w})$, where $k_{z}^{w}=-\arcsin(M/F_{z})$, and $k_{y}^{\pm}=\pi-\arccos[(\tilde{t}_{x}\pm\tilde{t}_{z}\cos k_{w}^{z})/\tilde{t}_{y}]$. By choosing the parameters such that $|\tilde{t}_{x}-\tilde{t}_{z}\cos k_{w}^{z}|<\tilde{t}_{y}<|\tilde{t}_{x}+\tilde{t}_{z}\cos k_{w}^{z}|$, only the two Weyl points at $\boldsymbol{k}_{w3}$ and $\boldsymbol{k}_{w4}$ are present. Correspondingly, there are two spin-down Weyl points located at $\boldsymbol{k}_{w3}$ and $\boldsymbol{k}_{w4}$. In Fig. 4(c), we plot the energy bands in the $k_{x}=0$ plane, which clearly exhibit two Weyl cones in each spin sector. Furthermore, Fig. 4(d) shows the distribution of Berry curvature vector field $\boldsymbol{\Omega}(\boldsymbol{k})=(\Omega_{x}(\boldsymbol{k}),\Omega_{y}(\boldsymbol{k}),\Omega_{z}(\boldsymbol{k}))$, where $\Omega_{a}(\boldsymbol{k})=\sum_{E_{n}<0}i\epsilon_{abc}\langle\partial_{k_{b}}u_{n}|\partial_{k_{c}}u_{n}\rangle$. Here $\epsilon_{abc}$ is the Levi-Civita symbol and $|u_{n}(\boldsymbol{k})\rangle$ is the eigenstate of the $n$-th occupied band. The field clearly shows that inversion-related Weyl points carry opposite monopole charges [103]. Specifically, a spin-down Weyl point at the top right (left) acts as a source (sink) of Berry flux, while its inversion partner—a spin-up Weyl point at the bottom left (right)—acts as a sink (source).
Discussions and conclusions.—We demonstrate that CPL can dynamically induce odd-parity altermagnetism in collinear $\mathcal{PT}$-symmetric antiferromagnets on dimerized lattices. The resulting $p$-wave altermagnetic phase inherently supports topological states due to the underlying Dirac band structure. Experimentally, the characteristic spin splitting can be detected by measuring the spin-split band structure along high symmetry momentum lines through spin- and angle-resolved photoemission spectroscopy.
@@
The antiferromagnets we consider are characterized by the coexistence of bond order and antiferromagnetic order. The bond order typically arises from electron-phonon interactions, whereas the antiferromagnetic order is driven by strong electron-electron repulsion. Their coexistence has been rigorously established through numerical studies of strongly correlated systems [96]. Beyond correlated materials, cold-atom systems offer a promising alternative platform for realizing our proposed odd-parity AMs. In such systems, the dimerized lattice can be directly engineered through optical lattice design [104], while the antiferromagnetic order can be stabilized by tuning the on-site interactions to be strongly repulsive [105, 106, 107]. The effect of CPL can be equivalently realized by periodically modulating the lattice [108].
As an outlook, odd-parity altermagnets also host favorable conditions for the emergence of unconventional superconductivity, as their band structure mimics time-reversal-invariant systems with SOC while explicitly breaking TRS. The fundamental relation $E(\mathbf{k},\boldsymbol{s})=E(-\mathbf{k},-\boldsymbol{s})$ ensures that a spin-up electron always has a spin-down partner available for pairing. This opens the door to realizing exotic, time-reversal-breaking spin-singlet superconductors.
@@
In conclusion, our work establishes a general dynamical protocol for creating odd-parity AMs, significantly expanding the class of systems capable of hosting these exotic phases and their emergent phenomena.